Test: Final Exam -(24 Questions This Question: 1 pt Consider these sample data 13, x25, x3 6, x4 19 a. Find n. b. Compute Σ C. Determine x. a. n c. x (Type an integer ar a decimal)

need answers asap please
Answer #1

a) here x1, x2,x3 and x4 4 samples are given so n= 4

b)x=x1+x2+x3 + 24-1345+6419 43

c) xbar =(x1+x2+x3+x4)/4 = 43/4 =10.75